OpenStudy (anonymous):

How do i find the derivative of (1+2cos(theta))(cos(theta)) ?

OpenStudy (anonymous):

deriv (1+2cos(t))*(cos(t)+(1+2cos(t)*deriv(cos(t)

Miracrown (miracrown):

deriv (1+2cos(t))*(cos(t))+(1+2cos(t)*deriv(cos(t) Was missing a paren. So, I think you're applying chain rule correctly

Miracrown (miracrown):

Two multiplicative terms ... derivative of them is (first * deriv of second) + (second * deriv of first) Do you know deriv of cos?

Miracrown (miracrown):

It is -sin. What about deriv (1 + 2cos(t))? What would that be?

Miracrown (miracrown):

Your participation is just as important as my help.

OpenStudy (anonymous):

Rusty on the rules, but i see it's the product rule here. However when I try to get the derivative of (1+2cost) I get (t-2sint), is this right? Because from here out it sways me away from the correct answer (which I already know, just not sure how to get there)

OpenStudy (anonymous):

sorry was on another tab trying to look for answers

Miracrown (miracrown):

That's alright. :) deriv of (1 + 2cost) = 0 - 2sint (deriv of 1 = 0, deriv of any constant = 0)

OpenStudy (anonymous):

Ha! Was thinking of integrals. Ok working on it now

OpenStudy (anonymous):

Got it :) Thanks a million!

Miracrown (miracrown):

Yw (: